A ten-year-old child is to receive 400 mg of acetaminophen every 4-6 hours as needed alternating with ibuprofen. Using acetaminophen liquid 160mg/5ml how much should the child receive per dose of acetaminophen?
Answer #1

ANSWER.

* Calculation of volume of acetaminophen receive by the child.

            Formula = Desired dose/Dose in hand * Quantity.

      Here

           Desired dose (D) = 400 mg.

           Dose in hand (H) = 160 mg.

                Quantity (Q) = 5 ml.

As per above formula

    Amount of acetaminophen receive per dose = D/H * Q

                                                                                 = 400/160 * 5

                                                                             = 2.5 * 5

                                                                                 = 12.5 ml.